# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v7

@@ 1.2 Related work and our contributions
Secondly, our study is conducted under a continuous spike-and-slab prior, to be introduced in Section 2. In contrast, Jiang [21], Jeong and Ghosal [20], and Tang and Martin [37] all study point-mass (discontinuous) spike-and-slab priors. In these other papers, a model complexity prior is used to first select a random subset of $s<n$ predictors. Conditionally on the chosen set, the $s$ coordinates are then endowed with a multivariate prior, while the other regression coefficients are modeled with a Dirac delta density at zero. In contrast, we employ an absolutely continuous prior which puts zero probability on exactly sparse vectors. Thus, our prior needs to be handled differently from these other papers.
Finally, we characterize the convergence rate of both a maximum a posteriori (MAP) estimator and the full posterior distribution. In contrast, Jiang [21], Jeong and Ghosal [20], and Tang and Martin [37] studied only the full posterior distribution but not any specific Bayesian point estimators. One may wonder why it is necessary to study MAP estimators and the posterior distribution separately. First, practitioners typically report a point estimate (e.g. the posterior mean, median, or mode) when they employ Bayesian methods. Thus, it is important to study the properties of these point estimators. Secondly, many researchers have shown that different Bayesian point estimates may have different asymptotic properties or behave very differently from the full posterior. For example, in the sparse normal means model, Johnstone and Silverman [22] showed that the posterior median under a point mass spike-and-slab prior attains the minimax risk, whereas the posterior mean converges at a slower, suboptimal rate. Under a different empirical Bayes spike-and-slab prior, Castillo and Mismer [10] showed that the posterior mean and median both obtain the optimal rate, but the full posterior converges at a suboptimal rate. In high-dimensional linear regression under a Laplace prior, Castillo et al. [11] showed that the posterior mode converges at the near minimax rate but the full posterior distribution converges much more slowly than the mode. These examples reinforce the argument that Bayesian point estimators need to be analyzed separately from the full posterior.
@@
The rest of this paper is structured as follows. In Section 2, we describe our prior specification and discuss Bayesian estimation of $\bm{\beta}$. In Section 3, we study the convergence rate of a local MAP estimator under our approach. In Section 4, we characterize the convergence rate for the entire posterior distribution and show that it can overcome the slow posterior contraction rate of the group lasso. Section 5 discusses how to implement our method. We conduct simulation studies in Section 6 and an analysis of an HIV drug resistance dataset in Section 7. Section 8 concludes the paper. Most of the proofs are deferred to the Appendix.
@@ 1.3 Notation and preliminaries
Note that $\mathbf{X}^{\top}\bm{\Omega}(\bm{\beta})\mathbf{X}$ is also the Fisher information matrix, or the expected value of the Fisher information matrix (1.6). Furthermore, if the canonical link function is used, then $\bm{\Sigma}(\bm{\beta}=\bm{\Omega}(\bm{\beta})$, and (1.7) can be greatly simplified to $\bm{\Sigma}(\bm{\beta})=\textrm{diag}\{b^{\prime\prime}(\mathbf{X}\bm{\beta})\}$.
@@ 2.2 Bayesian estimation
Local MAP estimators (or local posterior modes) are useful point estimates to consider because standard optimization algorithms can be used to rapidly obtain an estimate $\widehat{\bm{\beta}}$. These optimization algorithms are often faster and much more scalable than MCMC algorithms. In general, local MAP estimators may not be unique. However, as long as the prior $\pi(\bm{\beta})$ in (2.5) is proper and does not depend on the observed data, then the posterior density will also be proper and bounded, i.e. $\sup_{\bm{\beta}\in\mathbb{R}^{p}}\pi(\bm{\beta}\mid\mathbf{Y})<\infty$. This enables us to find a local MAP estimator (2.6).
@@ 3.1 Convergence rate of the MAP estimator
      For the function $\xi=(h\circ b^{\prime})^{-1}$ in (1.4), $\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta})<\infty$ for all $i=1,\ldots,n$.
@@
If we assume that the entries of $\mathbf{X}$ are uniformly bounded, i.e. $|x_{ij}|\leq D$ for some $D>0$, then we can relax Assumptions (A1) and (A2) so that $G$ and $s_{0}$ satisfy $\log G=o(n)$ and $s_{0}=o((n/\log G)^{1/2})$.
@@ 4.1 Posterior contraction rate
We first establish in Theorem 2 that the posterior $\pi(\bm{\beta}\mid\mathbf{Y})$ under the SSGL prior (2.2) asymptotically puts all of its mass on vectors where the generalized dimensionality (4.1) is no larger than a constant multiple of the true model size $s_{0}$. That is, the SSGL posterior concentrates on approximately sparse sets in high-dimensional GLMs. Theorem 3 then verifies that the SSGL posterior contracts at the minimax-optimal rate around the true parameter $\bm{\beta}_{0}$.
@@
$\mathbb{E}_{0}\Pi\left(\bm{\beta}:\lvert\bm{\nu}(\bm{\beta})\rvert>K_{1}s_{0}\mid\mathbf{Y}\right).\rightarrow 0$ (4.3)
@@ 4.2 Suboptimality of the group lasso for fully Bayesian inference
In Section 3.2, we demonstrated that there exists a local MAP estimator (2.6) for SSGL which converges at the minimax-optimal rate under weaker assumptions than those previously assumed for the group lasso (2.7). It turns out that from the fully Bayesian perspective, the SSGL also has an advantage over the group lasso. As discussed in Section 2.2, the group lasso estimator (2.7) is the MAP estimator under the prior $\pi(\bm{\beta})=\prod_{g=1}^{G}\bm{\Psi}(\bm{\beta}_{g}\mid\lambda)$, where $\bm{\Psi}(\cdot\mid\lambda)$ is a single multivariate Laplace density (2.1). In contrast to the two-group SSGL (2.2), however, the group lasso posterior might contract much slower than its posterior mode. This renders the group lasso less useful for uncertainty quantification. This is formalized in the next proposition.
@@
Suppose that we have a grouped GLM (3.1), and we endow $\bm{\beta}_{0}$ with the group lasso prior $\pi(\bm{\beta})=\prod_{g=1}^{G}\bm{\Psi}(\bm{\beta}_{g}\mid\lambda)$, where $\bm{\Psi}(\cdot\mid\lambda)$ is as in (2.1). Then there exist scenarios where the full posterior distribution $\pi(\bm{\beta}\mid\mathbf{Y})$ converges slower than the MAP estimator (2.7).
@@
Consider $\mathbf{Y}\sim\mathcal{N}_{n}(\bm{\beta}_{0},\bm{I}_{n})$, which falls under the exponential family (1.1) and is a special case of model (3.1), with $\mathbf{X}=\mathbf{I}_{n}$. First, suppose that $\bm{\beta}_{0}=\bm{0}_{n}$ and $m_{1}=\ldots=m_{G}=1$ so that the group lasso prior reduces to $\pi(\bm{\beta})=\prod_{g=1}^{G}(\lambda/2)\exp(-\lambda|\beta_{g}|)$. Setting $\lambda=\sqrt{2\log n}$ yields the near-minimax risk of $\sqrt{\log n}$ for the MAP estimator $\widehat{\bm{\beta}}$, i.e. $\lVert\widehat{\bm{\beta}}-\bm{\beta}_{0}\rVert_{2}=O_{p}(\sqrt{\log n})$ as $n\rightarrow\infty$ [12, 11]. However, by Theorem 7 of Castillo et al. [11], the posterior $\pi(\bm{\beta}\mid\mathbf{Y})$ then places no probability on the ball $\{\lVert\bm{\beta}\rVert_{2}\leq\sqrt{n/\log n}\}$ as $n\rightarrow\infty$. That is, for some $K_{4}>0$,
$\mathbb{E}_{\bm{\beta}_{0}=\bm{0}_{n}}\Pi\left(\bm{\beta}:\lVert\bm{\beta}\rVert_{2}\leq K_{4}\sqrt{\frac{n}{\log n}}~{}~{}\bigg{\lvert}~{}~{}\mathbf{Y}\right)\rightarrow 0~{}\textrm{ as }n\rightarrow\infty,$
and $\sqrt{\log n}\ll\sqrt{n/\log n}$ when $n$ is large. In the general group case where $\bm{\beta}_{0}=(\bm{\beta}_{01}^{\top},\ldots,\bm{\beta}_{0G})^{\top}$, suppose that $m_{g}=O(1)$ and $\bm{\beta}_{0g}=\bm{0}_{m_{g}}$ for all $g\in\{1,\ldots,G\}$. In this scenario, $\lambda=\sqrt{2\log n}$ still yields the near-minimax risk of $\sqrt{\log n}$ for $\widehat{\bm{\beta}}$ under $\ell_{2}$ loss. However, the posterior contraction rate under the group lasso prior $\pi(\bm{\beta})=\prod_{g=1}^{G}\bm{\Psi}(\bm{\beta}_{g}\mid\lambda)$ is still slower than $\sqrt{n/\log n}$.
Proposition 1 implies that the full posterior distribution under the group lasso prior may asymptotically put all of its mass in an $\ell_{2}$ ball with radius that is substantially larger than the convergence rate of the posterior mode. Intuitively, this is because there is a conflict between shrinking the coefficients to zero and optimally estimating the nonzero signals. In order to estimate very sparse parameters well, the group lasso needs to set $\lambda$ to be large. But if $\lambda$ is too large, then there will be too much bias in the resulting estimator. The addition of the slab density $\bm{\Psi}(\cdot\mid\lambda_{1})$ in the SSGL prior (2.2) alleviates this tension by preventing overshrinkage of true signals. Thus, the full SSGL posterior (2.4) also contracts at the minimax-optimal rate in $\ell_{2}$ error.
5 Implementation SSGL for GLMs
@@ 5.2 Choice of hyperparameters
The spike hyperparameter $\lambda_{0}$ in (5.4) controls how sparse our final model is, with larger values of $\lambda_{0}$ leading to more sparsity. For unstructured GLMs (1.2) with the SSL prior (2.3), Tang et al. [39] recommended tuning $\lambda_{0}$ from cross-validation (CV). For SSGL, we follow Tang et al. [39] and similarly tune $\lambda_{0}$ using $K$-fold CV, with a default of $K=10$. To accelerate the computational efficiency, we fit the model on each of the $K$ training sets in parallel, allowing for speed-ups on roughly the order of $K$. In practice, we tune $\lambda_{0}$ from an equispaced grid $\{\lambda_{0,1},\lambda_{0,2},\ldots,\lambda_{0,\max}\}$, where $0<\lambda_{0,1}<\lambda_{0,2}<\cdots<\lambda_{0,\max}$ and $\lambda_{0,\max}$ is the smallest value of $\lambda_{0}$ so that $\widehat{\bm{\beta}}=\bm{0}_{p}$. It is not hard to see that $\lambda_{\max}=\max_{1\leq g\leq G}\lVert\nabla_{g}~{}\ell_{n}(\bm{0}_{p})\rVert_{2}$, where $\nabla_{g}$ denotes the subvector of the gradient (1.5) corresponding to the $g$th group. Typically, $\lambda_{\max}$ has an analytical form. For example, in logistic regression, $\lambda_{\max}=\max_{1\leq g\leq G}\lVert 0.25\mathbf{X}_{g}^{\top}(\mathbf{Y}-0.5\bm{1}_{n})\rVert_{2}$.
@@ 5.3 Gibbs sampling for fully Bayesian inference
$\pi(\bm{\beta}\mid\bm{\omega},\mathbf{Y})$ $=\pi(\bm{\beta})\prod_{i=1}^{n}\mathcal{L}_{i}(\bm{\beta})\propto\pi(\bm{\beta})\exp\left\{-\frac{1}{2}(\mathbf{Z}-\mathbf{X}\bm{\beta})^{\top}\bm{\Omega}(\mathbf{Z}-\mathbf{X}\bm{\beta})\right\},$ (5.10)
where $\mathbf{Z}=(\kappa_{1}/\omega_{1},\ldots,\kappa_{n}/\omega_{n})^{\top}$, $\bm{\Omega}=\text{diag}(\omega_{1},\ldots,\omega_{n})$, and for $i=1,\ldots,n$, $\kappa_{i}=y_{i}-0.5$ and $\omega_{i}\sim\mathcal{PG}(1,0)$.
@@
$\pi(\bm{\beta}\mid\bm{\omega},\mathbf{Y})\approx\pi(\bm{\beta})\prod_{i=1}^{n}\widetilde{\mathcal{L}}_{i}(\bm{\beta})\overset{\cdot}{\propto}\pi(\bm{\beta})\exp\left\{-\frac{1}{2}(\widetilde{\mathbf{Z}}-\mathbf{X}\bm{\beta})^{\top}\bm{\Omega}(\widetilde{\mathbf{Z}}-\mathbf{X}\bm{\beta})\right\},$ (5.12)
where $\widetilde{\mathbf{Z}}=(\widetilde{\kappa}_{1}/\omega_{1}+\log M,\ldots,\widetilde{\kappa}_{n}/\omega_{n}+\log M)^{\top}$, $\bm{\Omega}=\text{diag}(\omega_{1},\ldots,\omega_{n})$, and for $i=1,\ldots,n$, $\widetilde{\kappa}_{i}=0.5(y_{i}-M)$ and $\omega_{i}\sim\mathcal{PG}(M,0)$.
@@
The conditional distribution (5.12) is not exact. However, (5.12) greatly simplifies posterior sampling for $\bm{\beta}$ in Poisson regression, since the updates for $\bm{\beta}$ can be obtained in closed form. Sampling from the exact conditional distribution of $\bm{\beta}$ in Poisson regression typically requires Metropolis-Hastings, which is not practical when $\bm{\beta}$ is high-dimensional. This is because there is a very low probability of accepting a proposal draw for $\bm{\beta}$ when $\bm{\beta}$ is a large vector. Therefore, we use the approximate conditional distribution (5.12). In simulation studies, we found that this approximation was very adequate.
@@
$\begin{array}[]{ll}\bm{\beta}\mid\bm{\tau}\sim\mathcal{N}(\bm{0},\bm{D}_{\bm{\tau}}),&\text{where }\bm{D}_{\bm{\tau}}=\text{Bdiag}(\tau_{1}\bm{I}_{m_{1}},\ldots,\tau_{G}\bm{I}_{m_{G}}),\\ \tau_{g}\mid\gamma_{g}\sim\text{Gamma}\left(\frac{m_{g}+1}{2},\frac{(\lambda_{g}^{\star})^{2}}{2}\right),&\text{where }\lambda_{g}^{\star}=\tau_{g}\lambda_{1}+(1-\tau_{g})\lambda_{0},\\ \gamma_{g}\mid\theta\sim\text{Bernoulli}(\theta),&g=1,\ldots,G.\end{array}$ (5.14)

### v8

@@ 1.2 Related work and our contributions
Secondly, our study is conducted under a continuous spike-and-slab prior, to be introduced in Section 2. In contrast, Jiang [22], Jeong and Ghosal [21], and Tang and Martin [40] all study point-mass (discontinuous) spike-and-slab priors. In these other papers, a model complexity prior is used to first select a random subset of $s<n$ predictors. Conditionally on the chosen set, the $s$ coordinates are then endowed with a multivariate prior, while the other regression coefficients are modeled with a Dirac delta density at zero. In contrast, we employ an absolutely continuous prior which needs to be handled differently from these other papers.
Finally, we characterize the convergence rate of both a maximum a posteriori (MAP) estimator and the full posterior distribution. In contrast, Jiang [22], Jeong and Ghosal [21], and Tang and Martin [40] studied only the full posterior distribution but not any specific Bayesian point estimators. One may wonder why it is necessary to study MAP estimators and the posterior distribution separately. First, practitioners typically report a point estimate (e.g. the posterior mean, median, or mode) when they employ Bayesian methods. Thus, it is important to study the properties of these point estimators. Secondly, many researchers have shown that different Bayesian point estimates may have different asymptotic properties or behave very differently from the full posterior. For example, in the sparse normal means model, Johnstone and Silverman [23] showed that the posterior median under a point mass spike-and-slab prior attains the minimax risk, whereas the posterior mean converges at a slower, suboptimal rate. Under a different empirical Bayes spike-and-slab prior, Castillo and Mismer [10] showed that the posterior mean and median both obtain the optimal rate, but the full posterior converges at a suboptimal rate. In high-dimensional linear regression under a Laplace prior, Castillo et al. [11] showed that the posterior mode converges at the near minimax rate but the full posterior distribution converges much more slowly than the mode. These examples reinforce the argument that Bayesian point estimators need to be analyzed separately from the full posterior. In this work, we focus on a local MAP estimator rather than the posterior mean or median, because of its practical appeal. Unlike the posterior mean or median, local MAP estimators under our spike-and-slab prior result in exact sparsity and are computationally faster to calculate. Given these appealing features, we aim to elucidate the theoretical optimality of MAP estimation for our model.
@@
The rest of this paper is structured as follows. In Section 2, we describe our prior specification and discuss Bayesian estimation of $\bm{\beta}$. In Section 3, we study the convergence rate of a local MAP estimator under our approach. In Section 4, we characterize the convergence rate for the entire posterior distribution and show that it can overcome the slow posterior contraction rate of the group lasso. Section 5 discusses how to implement our method. We conduct simulation studies in Section 6 and an analysis of an HIV drug resistance dataset in Section 7. Most of the proofs are deferred to the Appendix.
@@ 1.3 Notation and preliminaries
Note that $\mathbf{X}^{\top}\bm{\Omega}(\bm{\beta})\mathbf{X}$ is also the Fisher information matrix, or the expected value of the Fisher information matrix (1.6). Furthermore, if the canonical link function is used, then $\bm{\Sigma}(\bm{\beta})=\bm{\Omega}(\bm{\beta})$, and (1.7) can be greatly simplified to $\bm{\Sigma}(\bm{\beta})=\textrm{diag}\{b^{\prime\prime}(\mathbf{X}\bm{\beta})\}$.
@@ 2.2 Bayesian estimation
Local MAP estimators (or local posterior modes) are useful point estimates to consider because standard optimization algorithms can be used to rapidly obtain an estimate $\widehat{\bm{\beta}}$. These optimization algorithms are often faster and much more scalable than MCMC algorithms. In general, local MAP estimators may not be unique, and the likelihood function $\ell_{n}(\bm{\beta})$ may be unbounded. However, if the prior $\pi(\bm{\beta})$ in (2.5) is proper, does not depend on the observed data, and sufficiently restricts the parameter space for $\bm{\beta}$ (e.g. by shrinking most of the coefficients in $\bm{\beta}$ towards zero), then the posterior density $\pi(\bm{\beta}\mid\mathbf{Y})$ will be proper and bounded, i.e. ${\color[rgb]{0,0,0}\sup_{\bm{\beta}}}\pi(\bm{\beta}\mid\mathbf{Y})<\infty$. This enables us to find a local MAP estimator (2.6).
@@ 3.1 Convergence rate of the MAP estimator
      For the function $\xi=(h\circ b^{\prime})^{-1}$ in (1.4), $\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta})<\infty$ for all $\bm{\beta}\in\mathbb{R}^{p}$ and $i=1,\ldots,n$.
@@
By the continuous mapping theorem, $f(\widehat{\bm{\beta}})$ is also a consistent estimator of $f(\bm{\beta}_{0})$ for any function $f$ that is continuous at $\bm{\beta}_{0}$. For example, let $S\subset\{1,\ldots,p\}$, and let $\bm{\beta}_{S}$ be the subvector of $\bm{\beta}$ with indices in $S$. Taking $f(\bm{\beta})=\mathbf{A}_{S}\bm{\beta}$, where $\mathbf{A}_{S}\in\mathbb{R}^{|S|\times p}$ is a matrix with unit vector rows such that $\mathbf{A}_{S}\bm{\beta}=\bm{\beta}_{S}$, it follows that $\widehat{\bm{\beta}}_{S}$ is a consistent estimator of $\bm{\beta}_{S}$.
@@ 4.1 Posterior contraction rate
We first establish in Theorem 2 that the posterior $\pi(\bm{\beta}\mid\mathbf{Y})$ under the SSGL prior (2.2) asymptotically puts all of its mass on vectors where the generalized dimensionality (4.1) is no larger than a constant multiple of the true model size $s_{0}$. That is, the SSGL posterior concentrates on approximately sparse sets in high-dimensional GLMs. Theorem 3 then verifies that the SSGL posterior contracts at the minimax-optimal rate around the true parameter $\bm{\beta}_{0}$. In the subsequent theorems, the notation $\mathbb{E}_{0}$ denotes the expectation operator with the true parameter $\bm{\beta}_{0}$ in (3.1).
@@
$\mathbb{E}_{0}\Pi\left(\bm{\beta}:\lvert\bm{\nu}(\bm{\beta})\rvert>K_{1}s_{0}\mid\mathbf{Y}\right)\rightarrow 0$ (4.3)
@@
If $f(\bm{\beta})$ is a continuous function, then the posterior distribution for $f(\bm{\beta})$ is also consistent at $f(\bm{\beta}_{0})$. In particular, this implies that the marginal posteriors for the individual coefficients or subsets of the coefficients in $\bm{\beta}$ also concentrate around their respective true values (see Remark 2).
Remark 6.
@@ 4.2 Suboptimality of the group lasso for fully Bayesian inference
In Section 3.2, we demonstrated that there exists a local MAP estimator (2.6) for SSGL which converges at the minimax-optimal rate under weaker assumptions than those previously assumed for the group lasso (2.7). It turns out that from the fully Bayesian perspective, the SSGL also has an advantage over the group lasso. As discussed in Section 2.2, the group lasso estimator (2.7) is the MAP estimator under the prior $\pi(\bm{\beta})=\prod_{g=1}^{G}\bm{\Psi}(\bm{\beta}_{g}\mid\lambda)$, where $\bm{\Psi}(\cdot\mid\lambda)$ is a single multivariate Laplace density (2.1). In contrast to the two-group SSGL (2.2), however, the group lasso posterior might contract much slower than its posterior mode. This renders the group lasso less useful for uncertainty quantification. An example of this phenomenon is provided in the following proposition. Recall that $\bm{\beta}_{0}=(\bm{\beta}_{01}^{\top},\ldots,\bm{\beta}_{0G}^{\top})^{\top}$, where $\bm{\beta}_{0g}$ is the $g$th group of size $m_{g}$.
@@
Suppose that $\mathbf{Y}\sim\mathcal{N}_{n}(\bm{\beta}_{0},\mathbf{I}_{n})$ and all the group sizes are constant and equal, i.e. $m_{1}=\cdots=m_{G}=m$ where $1\leq m<\infty$. Further, assume that $\bm{\beta}_{01}\neq\bm{0}_{m}$ with $\lVert\bm{\beta}_{01}\rVert_{2}=L<\infty$, while $\bm{\beta}_{0g}=\bm{0}_{m}$ for all $g\in\{2,\ldots,G\}$. Suppose that we endow $\bm{\beta}_{0}$ with the prior $\pi(\bm{\beta})=\prod_{g=1}^{G}\bm{\Psi}(\bm{\beta}_{g}\mid\lambda)$, where $\bm{\Psi}(\cdot\mid\lambda)$ is the group lasso prior in (2.1). Then if $\lambda\asymp\sqrt{2\log n}$, the MAP estimator $\widehat{\bm{\beta}}$ satisfies
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}-\bm{\beta}_{0}\rVert_{2}\lesssim\sqrt{2\log n}\textrm{ as }n\rightarrow\infty,$ (4.6)
but for some $K_{4}>0$,
$\mathbb{E}_{0}\Pi\left(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq K_{4}\sqrt{\frac{n}{\log n}}~{}~{}\bigg{|}~{}~{}\mathbf{Y}\right)\rightarrow 0\textrm{ as }n\rightarrow\infty.$ (4.7)
@@
Appendix B. ∎
Proposition 1 implies that the full posterior distribution under the group lasso prior may asymptotically put all of its mass in an $\ell_{2}$ ball with radius that is substantially larger than the convergence rate of the posterior mode. Namely, under the grouped GLM (1.3) where $\mathbf{X}=\mathbf{I}_{n}$ and the exponential family (1.1) distribution is the normal distribution, the group lasso MAP estimator $\widehat{\bm{\beta}}$ has an expected $\ell_{2}$ risk of the order $\sqrt{2\log n}$ when the shrinkage parameter $\lambda$ is chosen of the order $\sqrt{2\log n}$. This is the minimax rate for estimating $\bm{\beta}_{0}$ [13]. However, with this same choice of $\lambda$, the posterior $\pi(\bm{\beta}\mid\mathbf{Y})$ places zero probability on the ball $\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq\sqrt{n/\log n}\}$ as $n\rightarrow\infty$, and $\sqrt{2\log n}\ll\sqrt{n/\log n}$ when $n$ is large. Intuitively, the discrepancy in the convergence rate between the MAP estimator and the full posterior is because there is a conflict between shrinking the coefficients to zero and optimally estimating the nonzero signals. In order to estimate very sparse parameters well, the group lasso needs to set $\lambda$ to be large; but if $\lambda$ is too large, then there is too much bias in the resulting estimator. The addition of the slab density $\bm{\Psi}(\cdot\mid\lambda_{1})$ in the SSGL prior (2.2) alleviates this tension by preventing overshrinkage of true signals. Thus, unlike the group lasso, the full SSGL posterior (2.4) also contracts at the minimax $\ell_{2}$ convergence rate.
5 Implementation of SSGL for GLMs
@@ 5.2 Choice of hyperparameters
The spike hyperparameter $\lambda_{0}$ in (5.4) controls how sparse our final model is, with larger values of $\lambda_{0}$ leading to more sparsity. For unstructured GLMs (1.2) with the SSL prior (2.3), Tang et al. [42] recommended tuning $\lambda_{0}$ from cross-validation (CV). For SSGL, we follow Tang et al. [42] and similarly tune $\lambda_{0}$ using $K$-fold CV, with a default of $K=10$. To accelerate the computational efficiency, we fit the model on each of the $K$ training sets in parallel, allowing for speed-ups on roughly the order of $K$. In practice, we tune $\lambda_{0}$ from an equispaced grid $\{\lambda_{0,1},\lambda_{0,2},\ldots,\lambda_{0,\max}\}$, where $0<\lambda_{0,1}<\lambda_{0,2}<\cdots<\lambda_{0,\max}$ and ${\color[rgb]{0,0,0}\lambda_{0,\max}}$ is the smallest value of $\lambda_{0}$ that results in a MAP estimator of $\widehat{\bm{\beta}}=\bm{0}_{p}$. It is not hard to see that ${\color[rgb]{0,0,0}\lambda_{0,\max}}=\max_{1\leq g\leq G}\lVert\nabla_{g}~{}\ell_{n}(\bm{0}_{p})\rVert_{2}$, where $\nabla_{g}$ denotes the subvector of the gradient (1.5) corresponding to the $g$th group. Typically, ${\color[rgb]{0,0,0}\lambda_{0,\max}}$ has an analytical form. For example, in logistic regression, ${\color[rgb]{0,0,0}\lambda_{0,\max}}=\max_{1\leq g\leq G}\lVert 0.25\mathbf{X}_{g}^{\top}(\mathbf{Y}-0.5\bm{1}_{n})\rVert_{2}$.
@@ 5.3 Gibbs sampling for fully Bayesian inference
$\pi(\bm{\beta}\mid\bm{\omega},\mathbf{Y})$ $=\pi(\bm{\beta})\prod_{i=1}^{n}\mathcal{L}_{i}(\bm{\beta})\propto\pi(\bm{\beta})\exp\left\{-\frac{1}{2}({\color[rgb]{0,0,0}\mathbf{Z}}-\mathbf{X}\bm{\beta})^{\top}\bm{\Omega}({\color[rgb]{0,0,0}\mathbf{Z}}-\mathbf{X}\bm{\beta})\right\},$ (5.10)
where ${\color[rgb]{0,0,0}\mathbf{Z}}=(\kappa_{1}/\omega_{1},\ldots,\kappa_{n}/\omega_{n})^{\top}$, $\bm{\Omega}=\text{diag}(\omega_{1},\ldots,\omega_{n})$, and for $i=1,\ldots,n$, $\kappa_{i}=y_{i}-0.5$ and $\omega_{i}\sim\mathcal{PG}(1,0)$.
@@
$\pi(\bm{\beta}\mid\bm{\omega},\mathbf{Y})\approx\pi(\bm{\beta})\prod_{i=1}^{n}\widetilde{\mathcal{L}}_{i}(\bm{\beta})\overset{\cdot}{\propto}\pi(\bm{\beta})\exp\left\{-\frac{1}{2}({\color[rgb]{0,0,0}\mathbf{Z}}-\mathbf{X}\bm{\beta})^{\top}\bm{\Omega}({\color[rgb]{0,0,0}\mathbf{Z}}-\mathbf{X}\bm{\beta})\right\},$ (5.12)
where ${\color[rgb]{0,0,0}\mathbf{Z}}=({\color[rgb]{0,0,0}\kappa_{1}}/\omega_{1}+\log M,\ldots,{\color[rgb]{0,0,0}\kappa_{n}}/\omega_{n}+\log M)^{\top}$, $\bm{\Omega}=\text{diag}(\omega_{1},\ldots,\omega_{n})$, and for $i=1,\ldots,n$, ${\color[rgb]{0,0,0}\kappa_{i}}=0.5(y_{i}-M)$ and $\omega_{i}\sim\mathcal{PG}(M,0)$.
@@
The conditional distribution (5.12) is not exact. However, (5.12) greatly simplifies posterior sampling for $\bm{\beta}$ in Poisson regression, since the updates for $\bm{\beta}$ can be obtained in closed form. Sampling from the exact conditional distribution of $\bm{\beta}$ in Poisson regression typically requires Metropolis-Hastings, which is not practical when $\bm{\beta}$ is high-dimensional. This is because we cannot decouple the individual components of $\bm{\beta}$ in the likelihood function, and therefore, we need to update the entire vector $\bm{\beta}$ in each MCMC iteration. When $\bm{\beta}$ is a large vector, there is a very low probability of accepting a proposal draw for $\bm{\beta}$, regardless of the proposal density used. Therefore, we sample $\bm{\beta}$ directly from the approximate conditional distribution (5.12). In simulation studies, we found that this approximation was very adequate.
@@
$\begin{array}[]{ll}\bm{\beta}\mid\bm{\tau}\sim\mathcal{N}_{p}(\bm{0}_{p},\bm{D}_{\bm{\tau}}),&\text{where }\bm{D}_{\bm{\tau}}=\text{Bdiag}(\tau_{1}\bm{I}_{m_{1}},\ldots,\tau_{G}\bm{I}_{m_{G}}),\\ \tau_{g}\mid\gamma_{g}\sim\text{Gamma}\left(\frac{m_{g}+1}{2},\frac{(\lambda_{g}^{\star})^{2}}{2}\right),&\text{where }{\color[rgb]{0,0,0}\lambda_{g}^{\star}=\gamma_{g}\lambda_{1}+(1-\gamma_{g})\lambda_{0}},\\ \gamma_{g}\mid\theta\sim\text{Bernoulli}(\theta),&g=1,\ldots,G.\end{array}$ (5.14)